Soybean meal is 12% protein; cornmeal is 6% protein. How many pounds of each should be mixed together in order to get 240-lb mixture that is 7% protein?
Answer by stanbon(57967)

Soybean meal is 12% protein; cornmeal is 6% protein. How many pounds of each should be mixed together in order to get 240-lb mixture that is 7% protein?
----------------
Let amt. of soy meal be "x" lbs. Amt. of protein is 0.12x lbs.
-------------------
Amt of corn meal is "240-x" lbs. Amt of protein is 0.06(240-x) = 14.4-0.6x lbs
-------------------
Amt of mixture is 240 lbs; Amt of protein is 0.07*240 = 16.8 lbs.
-----------------------
EQUATION:
protein + protein = protein
0.12x + 14.4-0.06x = 16.8
0.06x = 2.4
x = 40 lbs. (amount of soy meal in the mixture)
240-x = 200 lbs (amount of corn mean in the mixture)
---------------------------
Cheers,
Stan H.